# Charge carrier relaxation dynamics in the one-dimensional Kondo lattice model

### v1

@@ III.2 Dynamics in the presence of disorder
We next demonstrate that adding diagonal disorder suppresses the relaxation dynamics significantly [78, 79]. We first consider $J/t_{0}=0.5$ and then compute the decay of the spin polarization for increasing values of disorder strength $W$. Results obtained from averaging over at least 1216 disorder realizations are shown in Fig. 7(a) for $L=80$.
@@
Generically, we expect Anderson localization for a charge carrier in one dimension, yet coupling to a many-body bath may lift this [80, 81, 82]. Our data are consistent with localization on finite systems for sufficiently large $W/t_{0}$, as disorder detunes possible transitions via the coupling to the spin bath. In Fig. 7(b), we show the spin polarization in the large $J/t_{0}$ limit, where a simple picture emerges. The electron localizes in a site $i_{0}$ and the magnetization is coherently transferred back between the localized spin and the electronic spin. This picture is supported by the following argument: The oscillations at large $J/t_{0}$ are controlled by a frequency $\omega=J$, while the beating behavior observed without disorder is increasingly suppressed.
@@ III.3.1 Absence of thermalization in the $S^{z}_{\text{total}}=(L-1)/2$ sector
This behavior of $n_{k}$ is notably different from the one observed for the Holstein polaron in one dimension, where a thermal-like stationary $n_{k}$ with a broad momentum distribution centered around $k=0$ is quickly established for a comparable initial condition [21]. The Holstein-polaron chain, remarkably, exhibits clear eigenstate thermalization behavior already on small system sizes [47, 48]. However, for the Holstein polaron, the Hilbert-space scales exponentially with system size for the cases studied in [21, 47, 48].
@@
For our spin-polaron model and in the subspace with $S^{z}_{\text{total}}=(L-1)/2$, this is not the case. The simplest measure for quantum chaos is the gap ratio [83], defined as follows. Let $E_{n}$ be the many-body eigenenergies ordered as $E_{n+1}>E_{n}>E_{n-1}>\dots$, then the nearest-neighbor level spacing is defined as
@@
Note that we include 50% of all eigenstates centered around the mean energy in the averaging procedure. We find that the average gap ratio $r=\langle r_{n}\rangle$ in the $S^{z}_{\text{total}}=(L-1)/2$ is far away from GOE behavior expected for a quantum chaotic many-body system [63] (data not shown in the figures).
@@ IV.1 Two-electron case
In the first case, both electrons independently emit one magnon, exhibiting a dynamics that is overall similar to the one-electron case. Essentially, $n_{k}(t)$ exhibits two well-defined peaks at $k=-\pi/2+J/4$, and $k=\pi/2-J/4$. As time evolves, the function decays as an exponential function with a constant that is slightly renormalized from $J^{2}/4$. Note that in this four-body problem of two electrons and two magnons, interactions lead to a mid-spectrum bound state, studied in [57, 56].
In the second case, the up electron emits a magnon that can be reabsorbed by the down electron, leading to more interesting dynamics in the effective three-body problem. The spin-down electron momentum occupation at $k_{0}=\pi/2$ does essentially not decay, while the one of the spin-up electron decays. Notably, a small weight emerges at zero momentum, reminiscent of a bound state [57, 56, 55]. The exponential decay seen in the one-electron case gets more modified due the more complicated structure of decay channels in the two-particle case [56, 57]. In both cases, there is clearly no relaxation to a thermal state, as expected.
@@ IV.2 Towards a finite density of charge carriers
The same holds true for the decay of the total spin polarization of the conduction electrons, plotted in Fig. 11. $S_{c}^{z}(t)$ quickly decays and approaches a stationary value after a short time. The decay shows deviations from the exponential decay seen for one charge carrier due to the additional bands and the bound state in the spectrum [57, 56]. This stationary value is surprisingly close to the thermal expectation value (dashed line). To summarize, in this section, we numerically demonstrated that a system with just four electrons and four emitted magnetic excitations exhibits dynamics consistent with thermalization.
@@ V Decay of single charger carrier in a correlated and antiferromagnetic background
Figure 12: Average gap ratio $r$ of the model defined in Eq. (II.1) with $J_{\text{AFM}}/t_{0}=2$, evaluated in the $K=\pi/2$, $N_{e}=1$, and $S^{z}_{\text{total}}=(L-1)/2$ sector, as a function of the Kondo coupling strength for two different system sizes $L=12,16$. The spin exchanges of all spin-spin interactions contain an anisotropic spin exchange $\Delta=0.5$ [see Eq. (32)]. Here, we restrict our attention to the many-body eigenstates that are located in the middle half of the energy-ordered list of states for their sample, thereby avoiding high- and low-energy states. The dashed line represents the value of $r_{\text{GOE}}$ taken from [63].
@@
Figure 13: Relaxation dynamics in the Kondo-Heisenberg model. In (a), we show $n_{k}(t)$ as a function of quasi-momentum $k$ and time for the initial state from Eq. (33) with $L=16$, $J_{\text{AFM}}/t_{0}=2.0$, $J/t_{0}=2.0$, and $\Delta=0.5$. As time progresses, the quasi-momentum distribution develops a broad peak centered around $k=0$. (b) Time evolution of the conduction electron’s polarization $S^{z}_{c}$ (solid line) for the same initial state, compared to its thermal expectation value from the canonical ensemble (dashed lines) for $J_{\text{AFM}}/t_{0}=2.0$ and $\Delta=0.5$, and different values of $J/t_{0}$. We observe a reasonably good agreement at long times, given the small system sizes.
@@
In Fig. 13(b), we vary $J/t_{0}$ and plot the time evolution of $S_{c}^{z}$. While the relaxation dynamics clearly depend on $J/t_{0}$, for all selected parameters, the curves approach the thermal expectation value expected for the initial conditions. Obviously, due to the small system size, there are remaining oscillations, yet the overall trend is compatible with the expectations from the analysis of the gap ratio, that is, thermalization from the closed quantum system perspective [63].
@@ VI Summary and conclusion
We first provided a discussion of the well-studied case of one or two charge carriers in in a fully polarized spin background from the thermalization perspective. While exact results exist (see, e.g., [52]), a generic picture arises from considering the relevant electronic levels (in the case of one charge carrier, just two) coupled to a bosonic bath. Utilizing standard approaches, this generically leads to an exponential decay, in agreement with exact results [52] and with our exact numerical results. Furthermore, we investigated the effect of disorder on the relaxation dynamics, where the behavior at large values of disorder strength is consistent with localization. Most notably in the context of the current study, the expected absence of thermalization is evident in the stationary form of the quasi-momentum distribution for one or two charge carriers that exhibit resilient finite-momentum peaks.
@@
As an outlook, we envision to add optical phonons to the problem, which would allow us to study the competition between two bosonic baths. Technically, this could be simulated with advanced matrix-product state methods for phonons [84, 85, 86, 87] or a classical treatment of the phonons in the spirit of multitrajectory Ehrenfest dynamics [88] or the truncated Wigner approximation (see, e.g. [28]). Moreover, the initial state can be prepared as the result of an explicit modeling of an optical pulse.
Research data associated with the work is available at Zenodo [89].
@@ Acknowledgements.
We thank E. Jeckelmann, S. Kehrein, and S. Manmana for useful discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 217133147, 436382789; via CRC 1073 and large-equipment grants (GOEGrid).
@@ Appendix A Spectral decomposition of the initial state with one charge carrier in the fully polarized background
In Fig. 15, we observe a gap between ground state and excited states, which is proportional to the exchange coupling [32].
Figure 15: Energy spectrum as a function of the Kondo exchange coupling $J$ for a system size $L=216$ in the $K=\pi/2$ sector. The thin blue line is the ground-state energy, the thick orange line represents the band of excited states (see the inset for a zoom into that region). For any finite $J/t_{0}$, the band of excited states is separated from the ground state by an excitation gap of $\mathcal{O}(J)$. By calculating the spectral-weighted average of the local spin-spin operator from Eq. (34), we observe that the ground state is a singlet $K_{c,f}(\Psi_{\text{gs}})$] and the excited states are in the triplet sector $K_{c,f}(\Psi_{\text{ex}})$] for large $J/t_{0}$, as expected [32].
@@ Appendix B Two-level plus bath approximation
We briefly sketch the main approximations used in the analytical approach that leads to the exponential decay discussed in Sec. III.1.3, following [65] and [67]. The interaction between the two-level system and the bath is characterized by the coupling constant $\gamma_{q}$, with $q$ representing the wavevector of the emitted excitation in one dimension. In general, the tensor can be defined as:
@@
which tracks the transition probability between the two levels. A standard method to derive its dynamics is to use the Robertson projection operator technique [90, 91], which projects the full density matrix evolution onto the relevant subspace spanned by $\sigma_{e\rightarrow g}$. This results in an integro-differential equation involving memory kernels and the expectation value of $\sigma_{e\rightarrow g}$ at earlier times. However, solving this equation exactly is often intractable. To proceed, we assume Markovian dynamics (justified when the bath correlations decay rapidly), allowing us to neglect memory effects and replace the history-dependent term by its instantaneous value. This approximation simplifies the equation to a first-order differential equation. Finally, by taking the thermodynamic limit (replacing sums over $q$ by an integral), and introducing an exponential cutoff in time (see [65] for more detail), we obtain an exponential decay of the excited-state population, a hallmark of Markovian relaxation in open quantum systems.
@@
where the parameter $2{\gamma}$ characterizes the decay rate of the excited state in the Markovian limit. The function $f(\omega_{q})$ is a real and dimensionless quantity that expresses the frequency dependence of the magnon density states. To arrive at Eq. (43a) and Eq. (43b), we assumed the rotating-wave approximation (RWA) [92], which neglects rapidly oscillating counter-rotating terms. Now, at short times ($|\omega_{q}-\omega_{0}|t\ll 1$), if we consider $|\omega_{q}-\omega_{0}|>\gamma$, we can expand the exponential factor in Eq. (43a) in a Taylor series in $\gamma$. This yields an approximate expression for the amplitudes in order $\mathcal{O}(\gamma)$:

### v2

@@ III.2 Dynamics in the presence of disorder
We next demonstrate that adding diagonal disorder suppresses the relaxation dynamics significantly [nandkishore15, sierant25]. We first consider $J/t_{0}=0.5$ and then compute the decay of the spin polarization for increasing values of disorder strength $W$. Results obtained from averaging over at least 1216 disorder realizations are shown in Fig. 7(a) for $L=80$.
@@
Generically, we expect Anderson localization for a charge carrier in one dimension, yet coupling to a many-body bath may lift this [Bonca2018, Krause2021, Sierant2023]. Our data are consistent with localization on finite systems for sufficiently large $W/t_{0}$, as disorder detunes possible transitions via the coupling to the spin bath. In Fig. 7(b), we show the spin polarization in the large $J/t_{0}$ limit, where a simple picture emerges. The electron localizes in a site $i_{0}$ and the magnetization is coherently transferred back between the localized spin and the electronic spin. This picture is supported by the following argument: The oscillations at large $J/t_{0}$ are controlled by a frequency $\omega=J$, while the beating behavior observed without disorder is increasingly suppressed.
@@ III.3.1 Absence of thermalization in the $S^{z}_{\text{total}}=(L-1)/2$ sector
This behavior of $n_{k}$ is notably different from the one observed for the Holstein polaron in one dimension, where a thermal-like stationary $n_{k}$ with a broad momentum distribution centered around $k=0$ is quickly established for a comparable initial condition [Dorfner2015]. The Holstein-polaron chain, remarkably, exhibits clear eigenstate thermalization behavior already on small system sizes [Jansen2019, Schoenle2021]. However, for the Holstein polaron, the Hilbert-space scales exponentially with system size for the cases studied in [Dorfner2015, Jansen2019, Schoenle2021].
@@
For our spin-polaron model and in the subspace with $S^{z}_{\text{total}}=(L-1)/2$, this is not the case. The simplest measure for quantum chaos is the gap ratio [Oganesyan2007], defined as follows. Let $E_{n}$ be the many-body eigenenergies ordered as $E_{n+1}>E_{n}>E_{n-1}>\dots$, then the nearest-neighbor level spacing is defined as
@@
Note that we include 50% of all eigenstates centered around the mean energy in the averaging procedure. We find that the average gap ratio $r=\langle r_{n}\rangle$ in the $S^{z}_{\text{total}}=(L-1)/2$ is far away from GOE behavior expected for a quantum chaotic many-body system [DAlessio2016] (data not shown in the figures).
@@ IV.1 Two-electron case
In the first case, both electrons independently emit one magnon, exhibiting a dynamics that is overall similar to the one-electron case. Essentially, $n_{k}(t)$ exhibits two well-defined peaks at $k=-\pi/2+J/4$, and $k=\pi/2-J/4$. As time evolves, the function decays as an exponential function with a constant that is slightly renormalized from $J^{2}/4$. Note that in this four-body problem of two electrons and two magnons, interactions lead to a mid-spectrum bound state, studied in [Rausch2019, Moeller2012].
In the second case, the up electron emits a magnon that can be reabsorbed by the down electron, leading to more interesting dynamics in the effective three-body problem. The spin-down electron momentum occupation at $k_{0}=\pi/2$ does essentially not decay, while the one of the spin-up electron decays. Notably, a small weight emerges at zero momentum, reminiscent of a bound state [Rausch2019, Moeller2012, Moeller2012a]. The exponential decay seen in the one-electron case gets more modified due the more complicated structure of decay channels in the two-particle case [Moeller2012, Rausch2019]. In both cases, there is clearly no relaxation to a thermal state, as expected.
@@ IV.2 Towards a finite density of charge carriers
The same holds true for the decay of the total spin polarization of the conduction electrons, plotted in Fig. 11. $S_{c}^{z}(t)$ quickly decays and approaches a stationary value after a short time. The decay shows deviations from the exponential decay seen for one charge carrier due to the additional bands and the bound state in the spectrum [Rausch2019, Moeller2012]. This stationary value is surprisingly close to the thermal expectation value (dashed line). To summarize, in this section, we numerically demonstrated that a system with just four electrons and four emitted magnetic excitations exhibits dynamics consistent with thermalization.
@@ V Decay of single charger carrier in a correlated and antiferromagnetic background
Figure 12: Average gap ratio $r$ of the model defined in Eq. (II.1) with $J_{\text{AFM}}/t_{0}=2$, evaluated in the $K=\pi/2$, $N_{e}=1$, and $S^{z}_{\text{total}}=1/2$ sector, as a function of the Kondo coupling strength for two different system sizes $L=12,16$. The spin exchanges of all spin-spin interactions contain an anisotropic spin exchange $\Delta=0.5$ [see Eq. (32)]. Here, we restrict our attention to the many-body eigenstates that are located in the middle half of the energy-ordered list of states for their sample, thereby avoiding high- and low-energy states. The dashed line represents the value of $r_{\text{GOE}}$ taken from [DAlessio2016].
@@
Figure 13: Relaxation dynamics in the Kondo-Heisenberg model. In (a), we show $n_{k}(t)$ as a function of quasi-momentum $k$ and time for the initial state from Eq. (33) with $L=16$, $J_{\text{AFM}}/t_{0}=2.0$, $J/t_{0}=2.0$, and $\Delta=0.5$. As time progresses, the quasi-momentum distribution develops a broad peak centered around $k=0$. (b) Time evolution of the conduction electron’s polarization $S^{z}_{c}$ (solid line) for the same initial state, compared to its thermal expectation value from the canonical ensemble (dashed lines) for $J_{\text{AFM}}/t_{0}=2.0$ and $\Delta=0.5$, and different values of $J/t_{0}$. Specifically, for $J/t_{0}=1.0,\,2.0,$ and $4.0$, we evaluated the canonical ensemble expectation values at the temperatures consistent with the initial state, which yields $T=3.096\,t_{0}$, $3.112\,t_{0}$, and $3.375\,t_{0}$, respectively. We observe a reasonably good agreement at long times, given the small system sizes.
@@
In Fig. 13(b), we vary $J/t_{0}$ and plot the time evolution of $S_{c}^{z}$. While the relaxation dynamics clearly depend on $J/t_{0}$, for all selected parameters, the curves approach the thermal expectation value expected for the initial conditions. Obviously, due to the small system size, there are remaining oscillations, yet the overall trend is compatible with the expectations from the analysis of the gap ratio, that is, thermalization from the closed quantum system perspective [DAlessio2016].
@@ VI Summary and conclusion
We first provided a discussion of the well-studied case of one or two charge carriers in in a fully polarized spin background from the thermalization perspective. While exact results exist (see, e.g., [Henning2012]), a generic picture arises from considering the relevant electronic levels (in the case of one charge carrier, just two) coupled to a bosonic bath. Utilizing standard approaches, this generically leads to an exponential decay, in agreement with exact results [Henning2012] and with our exact numerical results. Furthermore, we investigated the effect of disorder on the relaxation dynamics, where the behavior at large values of disorder strength is consistent with localization. Most notably in the context of the current study, the expected absence of thermalization is evident in the stationary form of the quasi-momentum distribution for one or two charge carriers that exhibit resilient finite-momentum peaks.
@@
As an outlook, we envision to add optical phonons to the problem, which would allow us to study the competition between two bosonic baths. Technically, this could be simulated with advanced matrix-product state methods for phonons [Brockt2015, Jansen2020, Jansen2023, Koehler2021] or a classical treatment of the phonons in the spirit of multitrajectory Ehrenfest dynamics [tenBrink2022] or the truncated Wigner approximation (see, e.g. [Paprotzki2024]). Moreover, the initial state can be prepared as the result of an explicit modeling of an optical pulse.
Research data associated with the work is available at Zenodo [P-R_2025].
@@ Acknowledgements.
We thank A. Honecker, E. Jeckelmann, S. Kehrein, and S. Manmana for useful discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 217133147, 436382789, 493420525; via CRC 1073 and large-equipment grants (GOEGrid).
@@ Appendix A Spectral decomposition of the initial state with one charge carrier in the fully polarized background
In Fig. 15, we observe a gap between ground state and excited states, which is proportional to the exchange coupling [ueda1991].
Figure 15: Energy spectrum as a function of the Kondo exchange coupling $J$ for a system size $L=216$ in the $K=\pi/2$ sector. The thin blue line is the ground-state energy, the thick orange line represents the band of excited states (see the inset for a zoom into that region). For any finite $J/t_{0}$, the band of excited states is separated from the ground state by an excitation gap of $\mathcal{O}(J)$. By calculating the spectral-weighted average of the local spin-spin operator from Eq. (34), we observe that the ground state is a singlet $K_{c,f}(\Psi_{\text{gs}})$] and the excited states are in the triplet sector $K_{c,f}(\Psi_{\text{ex}})$] for large $J/t_{0}$, as expected [ueda1991].
@@ Appendix B Two-level plus bath approximation
We briefly sketch the main approximations used in the analytical approach that leads to the exponential decay discussed in Sec. III.1.3, following [seke1983] and [berman2010]. The interaction between the two-level system and the bath is characterized by the coupling constant $\gamma_{q}$, with $q$ representing the wavevector of the emitted excitation in one dimension. In general, the tensor can be defined as:
@@
which tracks the transition probability between the two levels. A standard method to derive its dynamics is to use the Robertson projection operator technique [robertson1966, seke1980], which projects the full density matrix evolution onto the relevant subspace spanned by $\sigma_{e\rightarrow g}$. This results in an integro-differential equation involving memory kernels and the expectation value of $\sigma_{e\rightarrow g}$ at earlier times. However, solving this equation exactly is often intractable. To proceed, we assume Markovian dynamics (justified when the bath correlations decay rapidly), allowing us to neglect memory effects and replace the history-dependent term by its instantaneous value. This approximation simplifies the equation to a first-order differential equation. Finally, by taking the thermodynamic limit (replacing sums over $q$ by an integral), and introducing an exponential cutoff in time (see [seke1983] for more detail), we obtain an exponential decay of the excited-state population, a hallmark of Markovian relaxation in open quantum systems.
@@
where the parameter $2{\gamma}$ characterizes the decay rate of the excited state in the Markovian limit. The function $f(\omega_{q})$ is a real and dimensionless quantity that expresses the frequency dependence of the magnon density states. To arrive at Eq. (43a) and Eq. (43b), we assumed the rotating-wave approximation (RWA) [cohen1992], which neglects rapidly oscillating counter-rotating terms. Now, at short times ($|\omega_{q}-\omega_{0}|t\ll 1$), if we consider $|\omega_{q}-\omega_{0}|>\gamma$, we can expand the exponential factor in Eq. (43a) in a Taylor series in $\gamma$. This yields an approximate expression for the amplitudes in order $\mathcal{O}(\gamma)$: